OpenStudy (anonymous):

suppose that $2000 is loaned at a rate of 18%, compounded monthly. Assuming that no payments are made, find the amount owed after 10 years. PLEASE HELP!!!!

ganeshie8 (ganeshie8):

its a straight forward problem. just plugin the numbers in "compound interest formula". may i know what exactly you need help wid ?

ganeshie8 (ganeshie8):

\(\huge FV = P(1 + \frac{r}{n})^{n*t}\) FV : Final Value of money P : Principal (initial money) r : rate of interest per year n : compounding period t : time in years

ganeshie8 (ganeshie8):

can you try plugin the given info in above formula and simplifying... .

OpenStudy (anonymous):

wht is compounded monthly mean?

ganeshie8 (ganeshie8):

it means, interest is calculated on loan every month. which means, 12 times a year each time at a rate of (r/12)

ganeshie8 (ganeshie8):

here, r = 18 %

OpenStudy (anonymous):

it means 1.5

ganeshie8 (ganeshie8):

thats right! it comes to 1.5 % every month. since there are 10 years, you have to compound this for 12*10 = 120 periods.

OpenStudy (anonymous):

u gave r and t as compounding pereid nd time r they mean same

ganeshie8 (ganeshie8):

nope. they are different. t = time is in years. (here t = 10 years) n = compoinding period, it refers to how many times you calculate interest in a year (here n = 12)

OpenStudy (anonymous):

kkk

ganeshie8 (ganeshie8):

all these are taken care by above formula, @emilybeth123 you plugin the numbers in the formula, you will be fine...